# The coupling mechanism between crossed-beams energy transfer and stimulated Brillouin scattering in homogeneous plasmas

### v1

@@ I Introduction
Controlling the laser plasma instabilities is an important task in inertial confinement fusion (ICF)ICF1 ; ICF3 ; D_D ; I_D ; ICFn , because the growth of these instabilities may affect the implosion symmetry of pellet compressingpress1 ; press2 . The crossed laser beams would induce the crossed-beams energy transfer when the phase match condition is satisfied. This instability is almost inevitable in ICF, however, people found that it is possible to tuning the implosion symmetry by CBETcbet1 ; cbet2 ; cbet3 .
@@ III The interaction between CBET and SBS
Figure 3: The theoretical CBET gain for different parameters. The results are obtained by Eq. (24). I stands for trivial regime of CBET, II stands for the depletion regime. (a) The CBET gain for different electron temperature and laser intensities when $\theta=31^{\circ}$. (b) The SBS gain for different electron temperature and laser intensities. III stands for the regime when the reflectivity is lower than $1\%$, IV stands for the regime when the reflectivity is higher than $1\%$,
@@
However, this equation dose not consider the influence of pump depletion, we should use Tang formulatang to obtain the final energy gain
@@
As an example, let’s consider two Gaussian laser with the same intensities, the width of laser is $90\lambda_{0}$ the crossed angle is set as $31^{\circ}$. The density of plasma is $0.05n_{c}$, the temperature of electron $T_{e}$ is varied, as shown in Fig 3, and the temperature of ions is $T_{i}=T_{e}/5$. The path length of SBS reflected light is $L/\rm{cos}(\theta/2)$, where $L=360\lambda_{0}$ is the length of plasma region. The seed level of SBS reflected light is $10^{-4}$.
@@ III.2 Fluid simulations for CBET and SBS interaction
Figure 4: The results of WEBS simulations. The intensities of laser is $1\times 10^{15}\rm{W/cm^{2}}$, the electron temperature is $4\rm{keV}$. (a) The time-dependent lasers’ energy for different cases. The red line and black line belong to the CBET only case, the blue dashed line obtained by CBET fluid theory, the red squares and black circles belong to the CBET and SBS case. (b) The time-dependent SBS reflectivity for lasers in different cases. The red line and black line belong to the SBS only case, and the red dashed line and black dashed line belong to the CBET and SBS case.
@@
Figure 7: The results of WEBS simulations. (a) The time-dependent lasers’ energy for different cases. The red line and black line belong to the CBET only case, the blue dashed line obtained by CBET fluid theory, the red dashed line and black dashed line belong to the CBET and SBS case. (b) The time-dependent SBS reflectivity for lasers in different cases. The red line and black line belong to the SBS only case, and the red dashed line and black dashed line belong to the CBET and SBS case.
@@
Figure 8: The simulation results by WEBS. (a) The laser intensities-dependent CBET gain, the orange diamonds are the results of CBET only case. the green dots stands for the results of CBET and SBS case. The black line is obtained by CBET fluid theory. (a) The plasma temperature-dependent CBET gain, the orange diamonds are the results of CBET only case. the green dots stands for the results of CBET and SBS case. The black line is obtained by CBET fluid theory.
@@ III.3 PIC simulations for CBET and SBS interaction with different $ZT_{e}/T_{i}$
In the PIC simulations, two Gaussian type laser beam enter into the simulation box with crossed angle $31^{\circ}$ ,the peak intensity og laser is $I_{0}=2.0\times 10^{15}\rm{W/cm^{2}}$ The SBS seed level is $5\times 10^{-4}I_{0}$. In Fig 10 (a), the electron temperature is $1\rm{keV}$, so, the parameter locateat the depletion regime of CBET which leads to strong interaction between CBET and SBS. As shown in Fig 10, the energy of Beam2 is transferred to SBS’s reflected light, the SBS reflectivity of Beam2 is much higher than Beam1. At around $20\rm{ps}$, we observe that the SBS was saturated, then the interaction between CBET and SBS becomes weaker. The reason of saturation is the strong laser filamentation instability(not show here).
@@ IV Conclusion and discussion
Figure 10: The PIC simulation results. (a) the time -dependent laser beam energy in CBET and SBS case.(b)the time -dependent SBS reflectivity in CBET and SBS case.
@@ V Acknowledgements
We are pleased to acknowledge useful discussions with Q. Wang, Y. G. Chen, S. Tan and W. B. Yao. This work was supported by the Scientific Research Foundation for High-level Talents of Anhui University of Science and Technology (Grant No.2022yjrc106), the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No.XDA25050700), National Natural Science Foundation of China (Grant Nos.11975059),and Natural Science Foundation of Hunan Province, China (Grant No.2020JJ5029).

### v2

@@ I Introduction
Controlling the laser plasma instabilities is an important task in inertial confinement fusion (ICF)ICF1 ; ICF3 ; D_D ; I_D ; ICFn , because the growth of these instabilities may affect the implosion symmetry of pellet compressingpress1 ; press2 . The crossed laser beams would induce the crossed-beams energy transfer (CBET) when the phase match condition is satisfied. This instability is almost inevitable in ICF, however, people found that it is possible to tuning the implosion symmetry by CBETcbet1 ; cbet2 ; cbet3 .
@@ III The interaction between CBET and SBS
Figure 3: The theoretical CBET gain for different parameters. The results are obtained by Eq. (24). (I) stands for trivial regime of CBET, (II) stands for the depletion regime. The $ZT_{e}/T_{i}$ is fixed at $10$. (a) The CBET gain for different electron temperature and laser intensities when $\theta=31^{\circ}$. (b) The SBS gain for different electron temperature and laser intensities. III stands for the regime when the reflectivity is lower than $1\%$, IV stands for the regime when the reflectivity is higher than $1\%$,
@@
However, this equation dose not consider the influence of pump depletion, we should use Tang formulatang to obtain the final spatial gain of CBET,
@@
As an example, let’s consider two Gaussian laser with the same intensities, the width of laser is $90\lambda_{0}$ the crossed angle is set as $31^{\circ}$. The density of plasma is $0.05n_{c}$, the temperature of electron $T_{e}$ is varied, as shown in Fig 3, and the temperature of ions is $T_{i}=T_{e}/5$. The path length of SBS reflected light is $L/\rm{cos}(\theta/2)$, where $L=360\lambda_{0}$ is the length of plasma region. The seed level of SBS reflected light is $\beta=10^{-4}$.
@@ III.2 Fluid simulations for CBET and SBS interaction
Figure 4: The results of WEBS simulations. The intensities of laser is $1\times 10^{15}\rm{W/cm^{2}}$, the electron temperature is $4\rm{keV}$,and the temperature of ions is $0.8\rm{keV}$. (a) The time-dependent lasers’ energy for different cases. The red line and black line belong to the CBET only case, the blue dashed line obtained by CBET fluid theory, the red squares and black circles belong to the CBET and SBS case. (b) The time-dependent SBS reflectivity for lasers in different cases. The red line and black line belong to the SBS only case, and the red dashed line and black dashed line belong to the CBET and SBS case.
@@
Figure 7: The results of WEBS simulations. The intensities of laser is $1\times 10^{15}\rm{W/cm^{2}}$, the electron temperature is $1\rm{keV}$,and the temperature of ions is $0.2\rm{keV}$. (a) The time-dependent lasers’ energy for different cases. The red line and black line belong to the CBET only case, the blue dashed line obtained by CBET fluid theory, the red dashed line and black dashed line belong to the CBET and SBS case. (b) The time-dependent SBS reflectivity for lasers in different cases. The red line and black line belong to the SBS only case, and the red dashed line and black dashed line belong to the CBET and SBS case.
@@
Figure 8: The simulation results by WEBS. (a) The laser intensities-dependent CBET gain, the orange diamonds are the results of CBET only case. the green dots stands for the results of CBET and SBS case. The black line is obtained by CBET fluid theory. The temperatures of electrons in (a) are $1\rm{keV}$,and the temperature of ions are $0.2\rm{keV}$. (b) The plasma temperature-dependent CBET gain, the orange diamonds are the results of CBET only case. the green dots stands for the results of CBET and SBS case. The black line is obtained by CBET fluid theory. The peak intensities of laser beams in (a) are $1\times 10^{15}\rm{W/cm^{2}}$.
@@ III.3 PIC simulations for CBET and SBS interaction with different $ZT_{e}/T_{i}$
Figure 10: The PIC simulation results. the laser intensities are $2\times 10^{15}\rm{W/cm^{2}}$, the electron temperatures are $1\rm{keV}$, the temperature of ions are varied in different cases. (a) the time-dependent laser beam energy in CBET and SBS case with $ZT_{e}/T_{i}=10$.(b)the time-dependent SBS reflectivity in CBET and SBS case with $ZT_{e}/T_{i}=10$.(c) the time-dependent laser beam energy in CBET and SBS case with $ZT_{e}/T_{i}=6$.(d)the time-dependent SBS reflectivity in CBET and SBS case with $ZT_{e}/T_{i}=6$. (e) the time-dependent laser beam energy in CBET and SBS case with $ZT_{e}/T_{i}=3$.(f)the time-dependent SBS reflectivity in CBET and SBS case with $ZT_{e}/T_{i}=3$.
@@
In the PIC simulations, two Gaussian type laser beam enter into the simulation box with crossed angle $31^{\circ}$ ,the peak intensity og laser is $I_{0}=2.0\times 10^{15}\rm{W/cm^{2}}$ The SBS seed level is $5\times 10^{-4}I_{0}$. In Fig 10 (a), the electron temperature is $1\rm{keV}$, so, the parameter locateat the depletion regime of CBET which leads to strong interaction between CBET and SBS. As shown in Fig 10, the energy of Beam2 is transferred to SBS’s reflected light, the SBS reflectivity of Beam2 is much higher than Beam1. At around $20\rm{ps}$, we observe that the SBS was saturated, then the interaction between CBET and SBS becomes weaker. The reason of saturation is the break up of IAWbreak , as show in Fig 11, the IAW of Beam2’s SBS starts to break up around the center of laser beam at $13.2\rm{ps}$, and the IAW already break up at $17.6\rm{ps}$.
Figure 11: The IAW of Beam2’s SBS in PIC simulations. (a) The IAW of Beam2’s SBS for $ZT_{e}/T_{i}=3$ case at $t=13.2$ $\rm{ps}$. (b) The IAW of Beam2’s SBS for $ZT_{e}/T_{i}=3$ case at $t=17.6$ $\rm{ps}$.
@@ V Acknowledgements
We are pleased to acknowledge useful discussions with Y. G. Chen, S. Tan and W. B. Yao. This work was supported by the Scientific Research Foundation for High-level Talents of Anhui University of Science and Technology (Grant No.2022yjrc106), the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No.XDA25050700), National Natural Science Foundation of China (Grant Nos.11975059),and Natural Science Foundation of Hunan Province, China (Grant No.2020JJ5029).